buds age is 1/7 the age of his uncle. if Bud is 8 years old,how old is his uncle
Answer by Maths68(1474)

Let x be the Uncle's age
Bud's age is 1/7 the age of his uncle
(1/7)x=8
Multiply by 7 both sides
7(1/7)x=8*7
(7/7)x=56
(1)x=56
x=56
Uncle is 56 years old